substitute equations
x=-4y-2
3x-y=20
what is the solution?

Answer by Jstrasner(112)

Hey,
So for this one we need to substitute a variable in one equation to what it equals in the other equation. In this case, we will substitute x in the second equation with -4y-2:
3(-4y-2)-y=20 => -12y-6-y=20 => -13y-6=20 => -13y=26 => -y=2 => y=-2
And then we substitute y in the first equation with -2:
x=-4(-2)-2 => x=8-2 => x=6
Therefore, x=6 and y=-2
I hope this helps!!